OpenStudy (anonymous):

Jim had 2.3 gallons of milk. If he drinks 1.2 quarts of milk everyday, how much milk will be there after 5 days. medal

OpenStudy (anonymous):

@Kinify

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

@alyssa_xo

OpenStudy (anonymous):

@Mehek14

Vocaloid (vocaloid):

@TINY10207: 1 gallon = 4 quarts 2.3 gallons = 2.3*4 = 9.2 quarts 5 days of drinking 1.2 quarts per day = 5*1.2 = 6 quarts gone 9.2 - 6 = 3.2 quarts left